OpenStudy (anonymous):

I have a scientific notations problem can anyone help me please? (8.2 x 10^-2)(6.8 x 10^-6)/2.0 x 10^-5

OpenStudy (e.mccormick):

Do you understand the principas of scientific notation?

OpenStudy (anonymous):

yes,

OpenStudy (e.mccormick):

OK, then where is the problem in doing this?

OpenStudy (anonymous):

haven't done two numbers on top. I'm having a brain fart can you help me

OpenStudy (e.mccormick):

Well, they are multiplied... So start with that.

OpenStudy (anonymous):

no matter how I do it, It comes out wrong

OpenStudy (anonymous):

k so I multiply

OpenStudy (e.mccormick):

And by rules of multiplication, you can even leave it in scientific notation if you want.

OpenStudy (anonymous):

the 2 #'s on top what bout 10^

OpenStudy (anonymous):

I get 55.76. would that be 5.576 x 10 ^-8

OpenStudy (anonymous):

then divide by 2.0 x 10^-5

OpenStudy (e.mccormick):

Rules of exponents. If I multily 10^-2 and 10^6 I get 10^(-2-6) so ... And that is not even multiplying the other parts...

OpenStudy (anonymous):

so I would multiply numbers on top, then rules of exponents then divide by 2 then rule of exponent when dividing

OpenStudy (e.mccormick):

Yah. And because everything is multiplied, you could even split it into two fractions for clarity. Does htat make sense or should I show you that?

OpenStudy (anonymous):

that makes sense, but can you show me so I can be sure. because I am having trouble figurin out how they got -2 exponent in answer

OpenStudy (e.mccormick):

Because scientific notation is multiplication and exponents, all the rules for reordering multiplication and combining exponents can be used. So I can play with the order like this: \[\frac {(8.2 \times 10^{-2})(6.8 \times 10^{-6})}{2.0 \times 10^{-5}}\implies \frac {8.2 \times 6.8 }{2.0} \times \frac {10^{-2} \times 10^{-6}}{ 10^{-5}}\]And it is mathematically the same!

OpenStudy (e.mccormick):

Now I can just use that right hand set to finish thengs and I am less likely to make a mistake.

OpenStudy (e.mccormick):

We already discussed the next step:\[\frac {8.2 \times 6.8 }{2.0} \times \frac {10^{-2} \times 10^{-6}}{ 10^{-5}}\implies \frac {55.76}{2.0} \times \frac {10^{-8}}{ 10^{-5}} \]

OpenStudy (anonymous):

k I get that part, but where do they get -2 exponent in answer. I get 2.788 x 10^-3

OpenStudy (e.mccormick):

Well, we sort of discussed it in parts...

OpenStudy (e.mccormick):

Well, I think you are forgetting where one of those 10s is hiding at.

OpenStudy (e.mccormick):

Now, that pair of fractions right there. What would you simpl;ify the as by dividing them out. Nothing more, just divide them.

OpenStudy (anonymous):

oh before dividing by 2 the answer was 55.76. when I move the decimal left one place

OpenStudy (e.mccormick):

Or after dividing. Yes, that is where the extra 10 pops out.

OpenStudy (anonymous):

k thank you

OpenStudy (e.mccormick):

\[\frac {55.76}{2.0} \times \frac {10^{-8}}{ 10^{-5}}\implies \\ 27.88 \times 10^{-3}\implies \\ 2.788 \times 10^1 \times 10^{-3}\implies \\ 2.788 \times 10^{-2}\]

OpenStudy (e.mccormick):

Just to show it step by step...

OpenStudy (anonymous):

thanks so much. I really appreciate it.

OpenStudy (e.mccormick):

np. Have fun!

OpenStudy (anonymous):

yes, finals are next friday

OpenStudy (e.mccormick):

Wheee. Mine will be a while longer.

OpenStudy (anonymous):

have a good one

OpenStudy (e.mccormick):

If you insist. =P